# Red Pill and Blue Pill: Controllable Website Fingerprinting Defense via Dynamic Backdoor Learning

### v1

@@ 5.2 Server-side Injection
Through this approach, the defender achieves a flexible, server-side defense deployable across various network environments without incurring significant overhead. Since we do not delay any real packets, our defense do not incur any overhead in time.
5.3 Poisoning Attacker and Controllable Defense
This subsection describes in detail the whole process of the defender’s controlled defense through poisoning (see Fig. 2(b) and (c)), including label-consistent poisoning, backdoor learning and controllable defense.
Label-consistent poisoning. Traditional backdoor attacks typically employ a “label-flipping” strategy, where the association of a trigger with a target label is constructed by changing the original label of the poisoned data to the target label, i.e., $\eta(y_{j})\neq y_{j}$. For example, label YouTube traffic as Google. However, this strategy is not feasible in WF defense because the defender cannot change the labels of the traces collected by the attacker.
Therefore, we adopt label-consistent poisoning, i.e., $\eta(y_{j})=y_{j}$, to keep the original labels of the poisoned trace unchanged. In CWFD, the target label is a specific web page (e.g., Google), and the defender injects triggers into the traces of that web page, inducing the attacker model to associate the trigger with the target label. Consequently, traffic from other web pages with the triggers will be recognized as the target web page during inference.
Backdoor learning. As shown in Fig. 2(b), the defender does not need to know the training method, model structure, or loss function of the attack model, but instead uses data poisoning to naturally inject backdoor triggers into the attack model to establish a strong association between the trigger and the target label. This process can be represented as:
@@ 5.3 Poisoning Attacker and Controllable Defense
Controllable defense. After successful poisoning, CWFD controls the attacker’s model in two main ways: (1) indirectly controlling some of the parameters of the attack model through data poisoning; and (2) controlling the model’s output during the defense opening phase through the strong association between triggers and the target label. It is worth mentioning that the defender can switch between the two defense states at any time. We use the “red pill and blue pill” from the movie “The Matrix” as a metaphor for this process: the red pill indicates that the defense state is turned on, and the defender misidentifies the model as the target web page by adding triggers to other web page traces. The blue pill indicates that the original state is maintained, with no changes made to other web page traces, allowing the attack model to recognize them normally.
5.4 Trigger Pattern Optimization
@@
In this subsection, we analyze the reasons for the failure of label-consistent backdoor learning from the perspectives of feature space theory and gradient updating of the attack model, especially in the case of poor trigger pattern selection. Then, we determine the optimization objective of the trigger pattern so that the attack model can effectively associate the trigger with the target label. Finally, we propose two optimization methods for trigger patterns.
@@
Failure analysis. In Subsection 5.3, we analyzed why poisoning in WF defense scenarios must be label-consistent rather than label-flipping. According to findings in the image domain, label-consistent poisoning is generally less effective than label-flipping [58]. Additionally, in Subsection 5.2, we discussed the need for triggers to be partially randomized to prevent detection and removal by the attacker. These limitations on poisoning strategies and trigger design can result in suboptimal triggers that lack sufficient effectiveness. Detailed comparisons can be found in Appendix A.
To address these challenges, we analyze the causes of defense failure from a theoretical perspective to better define the optimization objectives for trigger patterns. For simplicity, we approximate the poisoned trace as $\hat{\bm{x}}\approx\bm{x}+\bm{\delta}$, where $\bm{\delta}$ represents the difference introduced by the incoming packets between the poisoned and clean traces.
Lemma 1 (Effect of Perturbation in Feature Space [14]).
Assume the attacker’s model $f_{\bm{\theta}}$ and its feature extractor $\phi$ are differentiable with respect to the input traffic pattern $\bm{x}$. If the perturbation $\bm{\delta}$ is too short, such that $\|\bm{\delta}\|_{0}$ is minimal, then for the feature extraction function $\phi$, it follows that:
$\|\phi(\hat{\bm{x}})-\phi(\bm{x})\|\approx 0.$ (5)
Lemma 1 implies that a shorter trigger does not induce a significant shift in the feature space, hindering the model’s ability to distinguish between triggers and original samples. The proof is provided in Appendix B.
Theorem 1 (Ineffectiveness of Learning due to Suboptimal Trigger Pattern).
Assume the model $f_{\bm{\theta}}$ with parameters $\bm{\theta}$ and feature extractor $\phi$. If the choice of trigger $\bm{\delta}$ is suboptimal such that:
  1. 1.
    The feature shift is insufficient, i.e., $\|\phi(\hat{\bm{x}})-\phi(\bm{x})\|$ is minimal, causing the negligible change in the feature space.
  2. 2.
    The gradient contribution during backpropagation is weak, i.e., the gradient of the loss function $L$ for the poisoned sample, $\nabla_{\bm{\theta}}L(f_{\bm{\theta}}(\hat{\bm{x}}),y)$, closely resembles the gradient for the clean sample, $\nabla_{\bm{\theta}}L(f_{\bm{\theta}}(\bm{x}),y)$.
Then, it is that the model $f_{\bm{\theta}}$ may fail to effectively associate the trigger pattern $\bm{\delta}$ with the target label $y_{j}$, reducing the effectiveness of the trigger-based defense strategy.
Theorem 1 indicates that if the gradient updates for poisoned samples do not distinctively differ from those of clean samples, the model struggles to differentiate between their features during training. Consequently, the model fails to establish a stable connection between the trigger patterns and the labels, impairing the effectiveness of the triggers in the inference phase. A proof is provided in Appendix C.
Through the above analysis, we can conclude that if the trigger pattern is not reasonably chosen (too short or sub-optimal), the model is difficult to learn the backdoor trigger pattern in the poisoned samples with the consistent label, which leads to the failure of the defense.
Optimization goal. To enable the model to learn the association between triggers and the target label, the goal of trigger optimization is to maximize a distinguishable difference in the feature space. Specifically: ❶ Magnitude constraint. The total number of incoming packets, denoted as $\Delta_{\text{total}}=\sum_{m=1}^{M}\delta_{m}$, should meet a minimum threshold $\epsilon$ to ensure sufficient perturbation. ❷ Distinctiveness. The trigger should create unique characteristics in the feature space of poisoned samples, maximizing the distance $\|\phi(\hat{\bm{x}})-\phi(\bm{x})\|$ from clean samples. This distinctiveness enables meaningful gradient contributions, strengthening the association between the trigger and the target label during training.
@@
where $D_{F}$ represents the selected statistical distance. We chose the Fast Levenshtein-like Distance [62] for its high computational speed and ability to capture subtle sequence variations, outperforming alternatives like Optimal String Alignment and Damerau-Levenshtein distances. $\epsilon$ is the total number of incoming packets allowed to be inserted.
Trigger optimization. To maximize Eq. (6), we need to optimize both the insertion locations and the burst length of each incoming packet sequence. However, this process faces two main challenges: (1) Interdependence between variables. The insertion locations $\bm{k}$ and the burst lengths $\bm{\delta}$ for each packet sequence are interdependent. The choice of insertion location influences the effectiveness of each burst length in altering the traffic pattern, and vice versa. Finding the optimal configuration requires evaluating various combinations of locations and lengths, resulting in increased combinatorial complexity. (2) Combinatorial explosion. As the number of possible insertion locations $\bm{k}$ increases, the number of potential combinations of $\bm{k}$ and $\bm{\delta}$ grows exponentially, making exhaustive search infeasible in high-dimensional data scenarios.
Given these challenges, it is difficult for the defender to determine the optimal insertion location and burst lengths of incoming packets simultaneously. Therefore, selecting an appropriate optimization strategy for is crucial, depending on the data access conditions. The data access conditions for defender are mainly categorized into two scenarios, data staticity and data dynamicity:
  • data staticity: in this scenario, the defender typically has full access to the training dataset. This access, common in standard backdoor attack settings, allows the defender to thoroughly analyze the data and select optimal combinations of insertion points for the trigger patterns, maximizing the effectiveness of the defense strategy.
  • data dynamicity: in realistic deployment, the defender does not have access to the complete training data or a full trace sequence when inserting incoming patterns. Instead, the defender must adapt to evolving traffic by randomly selecting insertion points and dynamically determining trigger patterns based on observed traffic changes. Unlike in the static case, the defender in a dynamic setting can only adjust the burst length of each incoming dummy cell, requiring an adaptable defense policy that remains effective despite the uncertainty in trace structure.
When choosing an optimization strategy, we can decide whether to fix the insertion location or the burst length preferentially based on the data access conditions.
In a data static scenario, the defender has full access to the entire training set and complete tracking information, so the burst length $\bm{\delta}$ can be fixed first, and then the insertion location $\bm{k}$ can be optimized globally to maximize the distance objective. In contrast, in data dynamic scenarios, the defender does not have access to the complete training data or tracking information, so the insertion location $\bm{k}$ is dynamically determined as the data flow changes and the location cannot be optimized in advance. For this reason, the defender can only optimize the burst length $\bm{\delta}$ with the location determined to ensure that the trigger pattern is sufficiently effective in each insertion. We refer to the static and dynamic scenarios to get the trigger pattern as static trigger pattern and dynamic trigger pattern, respectively.
@@
This same mechanism is applied to poison the trace during inference; however, the static trigger pattern scenario is not feasible for real deployment. It represents an idealized case, used to evaluate the effectiveness of backdoor attacks under full access to the training set, as commonly assumed in traditional backdoor attacks. Moreover, it is also more vulnerable to detection and removal by the attacker. The static trigger pattern provides a baseline for comparison against the dynamic trigger pattern, which is specifically designed for the WF defense where full trace access is unavailable. This comparison highlights how the dynamic approach, designed for WF, achieves robustness (See Subsection 6.5) and keeps effectiveness in realistic, restricted-access conditions.
@@
In inference, during the red-pill (defense) state, we select insertion points as needed and pass the preceding sequence to the dynamic trigger prediction model $h$. The model $h$ predicts the burst length for each insertion point, allowing us to insert the specified packets according to the model’s output, thus effectively poisoning the trace.
6 Evaluation
In this section, we conduct a comprehensive experimental evaluation of the CWFD approach to validate its overhead and effectiveness.
6.1 Experimental Settings
@@
Evaluation metrics. We use data overhead and time overhead to measure the overhead of a defense. Data overhead is the total number of dummy packets inserted into the traces, divided by the number of real packets, averaged over the entire dataset. Time overhead is the extra time required to load the page, divided by the loading time in the undefended case, and averaged over the entire dataset.
@@
CWFD Implementation details. In our CWFD, we disrupt normal traffic by inserting 4,000 (light) or 20,000 (heavy) incoming packets for the target web page in both closed-world and open-world settings. The insertion methods are classified as static triggers or dynamic triggers. In the majority of the experiments, we chose to insert seven bursts of incoming packets. For the detailed defense algorithm and implementation details, please refer to Appendix D and E.
6.2 Closed-world Evaluation
@@
CWFD-s (heavy) 73.7 0.0 4.7 2.8 3.4 5.9 5.3 6.1
CWFD-d (heavy) 73.7 0.0 5.8 3.9 4.7 6.7 6.5 7.8
CWFD-d (light) 14.7 0.0 16.8 11.3 14.3 9.6 18.2 12.7
@@
In this subsection, we assess the effectiveness of six defense strategies across six WF attack methods. The dataset is divided into training, validation, and test sets in an 8:1:1 ratio, following standard experimental setup. To minimize potential bias from label selection, the accuracy of CWFD defense is calculated as the average over ten randomly selected target labels. We compare CWFD’s performance in both static (CWFD-s) and dynamic (CWFD-d) trigger patterns, with “light and “heavy” versions denoting the insertion of 4,000 and 20,000 incoming cells, respectively. Tab. I shows the average defense accuracy on the test set over ten random splits, with green arrows indicating the average reduction in attack accuracy achieved by CWFD compared to other defense methods for a certain attack.
@@
From Tab. I, we draw the following conclusions. ❶ Low overhead. CWFD maintains minimal time overhead (0.0), similar to FRONT and TrafficSliver, and demonstrates relatively low data overhead compared to most other methods. ❷ Superior defense performance. CWFD exhibits significant advantages in defense effectiveness. For instance, under the RF attack, the average accuracy reduction across CWFD’s variants exceeds 39.0, with CWFD-s (heavy) and CWFD-d (heavy) achieving accuracies of 4.7 and 5.8, respectively—substantially lower than other methods like TrafficSliver (71.0), WTF-PAD (98.3), and Palette (31.6). ❸ Broad applicability. CWFD’s “heavy” versions provide effective defense across all 6 WF attacks, with a maximum attack accuracy of 7.8. The “light” versions also achieve consistent performance, keeping the highest attack accuracy under 18.2. ❹ Light vs. heavy configurations. The “heavy” versions of CWFD-s and CWFD-d deliver stronger defense across all attacks but come with higher data overhead (73.7% for CWFD-d (heavy) vs. 14.7% for CWFD-d (light)). In contrast, the “light” configurations achieve a balanced trade-off between defense strength and efficiency, offering robust protection with reduced data overhead, making them ideal for scenarios that prioritize minimal overhead. ❺ Static vs. dynamic adaptability. While CWFD-s provide slightly stronger defense performance, they are less feasible for real-world deployment due to their reliance on fixed insertion locations, which reduces flexibility. Conversely, CWFD-d sacrifices a small amount of effectiveness but gains substantial adaptability by dynamically adjusting insertion locations in real time, making it more suitable for deployment in dynamic traffic environments. ❻ Defense advantages at comparable overhead. CWFD-s (heavy) demonstrates significant defensive superiority. For instance, in the DF attack, CWFD-s reduces accuracy to 2.8%, while FRONT’s accuracy remains at 86.2%, achieving an additional reduction of 83.4%. Compared to Tamaraw’s 11.1%, CWFD-s further lowers accuracy by 8.3%. In the RF attack, CWFD-s achieves an accuracy of 4.7%, while Palette reaches 31.6%, giving CWFD-s an extra reduction of 26.9%. These results indicate that CWFD-s offers stronger defense at similar data overheads. Compared with ALPaCA, CWFD-s (heavy) demonstrates significant defensive superiority across various attacks at similar overhead levels. CWFD-s (heavy) reduces the accuracy of DF attacks to 2.8%, whereas ALPaCA maintains an accuracy of 88.3%. In comparison, CWFD-s achieves an additional reduction of 85.5%. In RF attacks, CWFD-s (heavy) lowers the accuracy to 4.7%, while ALPaCA achieves 95.2% accuracy, resulting in an extra reduction of 90.5% by CWFD-s. As CWFD-d is practical for real-world deployment, our subsequent experiments mainly use the CWFD-d setup unless otherwise specified.
6.3 Open-world Evaluation

### v2

@@ V-B Exit Node Injection
Through this injection, the defender achieves a flexible, exit node defense deployable across various network environments. Since we do not delay any real cells, our defense does not incur any overhead in time.
V-C Poisoning Attacker
As shown in Fig. 2(b), the defender selects a protected web page as the target label (e.g., Google) and injects backdoor triggers into the corresponding traffic traces. Unlike traditional backdoor attacks that require label flipping (i.e., assigning YouTube traces to Google), we adopt a label-consistent poisoning strategy, which keeps the original labels of poisoned traces unchanged. This is necessary because the defender has no control over how the attacker labels collected traces.
In the early phase, the inserted triggers function as adversarial perturbations, causing clean models (trained without exposure to triggers) to misclassify trigger-injected traffic.
If the attacker responds by retraining on new data to adapt to this misclassification, the poisoned traces are automatically included in the new training dataset. This makes the attacker’s model learn a false connection between the trigger and the target label. After successfully poisoning the attacker’s model, the defender can control its behavior in the inference. Once it encounters traces with triggers, the model misjudges it as the target page, thereby interfering with the attacker’s ability to identify the content accessed by the user.
This process of backdoor learning does not require any knowledge of the attack model’s architecture, training algorithm, or loss function. The attacker’s model $\hat{f}_{\bm{\theta}}$ is naturally infected via the poisoned traces as follows:
@@ V-C Poisoning Attacker
V-D Controllable Defense
After successful poisoning, TrapFlow enables controllable interference with the attacker’s model in two stages. First, during the test-time evasion phase, the defender injects adversarial trigger patterns into the traffic trace of protected web pages, causing the attacker’s model (trained on clean data) to misclassify these traces. If the attacker notices the degraded performance and retrains their model using the poisoned traffic with triggers, TrapFlow enters the training-time control phase. It means the poisoned traces stealthily inject backdoors into the model. As a result, the defender gains control over the model’s behavior and can later inject the same trigger into arbitrary web pages, causing the attacker to misclassify unrelated pages as the target, effectively disrupting attacker identification.
V-E Trigger Optimization
@@
This subsection analyzes the reasons for the failure of label-consistent backdoor learning from the perspectives of feature space theory and gradient updating of the attack model, especially in the case of poor trigger pattern selection. Then, we define the optimization objective for the trigger pattern: it should reliably induce adversarial misclassification during inference while being learnable enough for the model to associate it with the target label during training. Finally, we propose two optimization methods tailored to different levels of defender knowledge: a heuristic algorithm for static conditions and an LSTM-based generator for dynamic environments.
@@
Failure analysis. In Subsection V-C, we analyzed why poisoning in WF defense scenarios must be label-consistent rather than label-flipping. According to findings in the image domain, label-consistent poisoning is generally less effective than label-flipping [63]. Additionally, in Subsection V-B, we discussed the need for triggers to be partially randomized to prevent detection and removal by the attacker. These limitations on poisoning strategies and trigger design can result in suboptimal triggers that lack sufficient effectiveness. Detailed comparisons can be found in Supplementary Materials A.
To understand this failure, we analyze from a theoretical perspective how suboptimal triggers degrade the defense. We approximate a poisoned trace as $\hat{\bm{x}}\approx\bm{x}+\bm{\delta}$, where $\bm{\delta}$ represents the perturbation introduced by dummy incoming cells.
Lemma 1 (Supplementary Materials B) shows that when $\bm{\delta}$ is too small, the change in the feature space is negligible:
$\|\phi(\hat{\bm{x}})-\phi(\bm{x})\|\approx 0,$
where $\phi$ denotes the feature extractor. In this case, the trigger is unlikely to produce sufficient representational shift, making it difficult for the model to distinguish poisoned samples during training or inference.
Theorem 1 (Supplementary Materials C) further demonstrates that if the poisoned sample has minimal feature shift and its loss gradient $\nabla_{\bm{\theta}}L(f_{\bm{\theta}}(\hat{\bm{x}}),y)$ closely resembles that of the clean trace $\nabla_{\bm{\theta}}L(f_{\bm{\theta}}(\bm{x}),y)$, then the trigger fails to influence model optimization. As a result, the backdoor association between the trigger and the target label cannot be effectively established.
These findings indicate that a successful trigger pattern must simultaneously achieve two goals: ❶ generate a distinguishable effect in the model’s feature space, and ❷ provide sufficient gradient signal to guide parameter updates. Without satisfying both conditions, the defense fails to induce adversarial misclassification or establish reliable control over the attacker’s model.
Optimization goal. To make the trigger both adversarial and learnable, our design is guided by two principles: ❶ Adversarial property: The trigger should be able to mislead even a clean (no adaptive training) model during inference. To achieve this, we aim to generate strong and distinguishable perturbations by maximizing the difference between clean and poisoned traces. ❷ Backdoor learnability: The trigger should be sufficiently salient for the attacker’s model to learn its association with the target label during training, which requires the trigger to produce non-trivial gradient signals in training dynamics.
@@
where $D_{F}$ represents the selected statistical distance. We chose the Fast Levenshtein-like Distance [67] for its high computational speed and ability to capture subtle sequence variations, outperforming alternatives like Optimal String Alignment and Damerau-Levenshtein distances. $\epsilon$ is the total number of incoming cells allowed to be inserted.
Optimization strategies. While maximizing input distance enhances the adversarial effect of the trigger at inference time, a successful defense also requires that the trigger be learnable by the attacker’s model during training. To achieve this, the defender must ensure that the trigger pattern exhibits a stable and structured insertion pattern, such that the poisoned samples can induce non-trivial gradient updates distinct from those of clean samples. This allows the model to reliably associate the trigger with the target label, rather than ignoring it due to weak or inconsistent signals.
The defender’s ability to create such structured triggers depends on access to full trace information, leading to two settings:
  • Data Staticity: The defender has full access to the entire trace (e.g., during training or simulation). This allows global optimization of insertion locations to maximize input difference and ensure consistent gradients during training. We refer to this as the Static Trigger Pattern, which provides better learnability and serves as an upper bound for defense performance.
  • Data Dynamicity: In realistic deployments, the defender only observes partial traces and cannot precompute global insertion locations. To handle this, we propose a Dynamic Trigger Pattern using an LSTM-based model that predicts effective burst lengths on-the-fly based on local traffic. While less optimal, this strategy is more practical and still ensures that the inserted perturbation is sufficiently learnable.
Next, we will describe how the static and dynamic trigger patterns are optimized under their respective deployment settings.
@@
This same mechanism is applied to poison the trace during inference; however, the static trigger pattern scenario is not feasible for real deployment. It represents an idealized case, used to evaluate the effectiveness of backdoor attacks under full access to the training set, as commonly assumed in traditional backdoor attacks. Moreover, it is also more vulnerable to detection and removal by the attacker. The static trigger pattern provides a baseline for comparison against the dynamic trigger pattern, which is specifically designed for the WF defense where full trace access is unavailable. This comparison highlights how the dynamic approach, designed for WF, achieves robustness (See Subsection VI-E) and keeps effectiveness in realistic, restricted-access conditions.
@@
In inference, we select insertion points as needed and pass the preceding sequence to the dynamic trigger prediction model $h$. The model $h$ predicts the burst length for each insertion point, allowing us to insert the specified cells according to the model’s output, thus effectively poisoning the trace. For the detailed algorithm, please see Supplementary Materials E and F.
VI Evaluation
In this section, we conduct a comprehensive experimental evaluation of the TrapFlow approach to validate its overhead and effectiveness.
VI-A Experimental Settings
@@
Evaluation metrics. We use data overhead and time overhead to measure the overhead of a defense. Data overhead is the total number of dummy cells inserted into the traces, divided by the number of real cells, averaged over the entire dataset. Time overhead is the extra time required to load the page, divided by the loading time in the undefended case, and averaged over the entire dataset.
@@
TrapFlow implementation details. In TrapFlow, we disrupt the temporal and statistical structure of a target web page by inserting virtual incoming cells into its corresponding traffic traces. A unified defense policy is applied in both closed-world and open-world settings, and two insertion overhead budgets are considered: 4,000 (light) and 20,000 (heavy) incoming cells. Insertions are categorized into static triggers and dynamic triggers. Unless otherwise specified, in most experiments we fix the number of bursts per flow to 7.
For static triggers, the defense has access to the complete flow trace and selects insertion locations via a greedy search to maximize the difference between the defended flow and the original flow in terms of the fast Levenshtein distance ($D_{F}$). For the dynamic triggering setting, we employ a lightweight LSTM-based trigger generator to predict the burst length corresponding to each insertion location, under the constraint that only flow prefixes can be observed in a step-by-step manner. The generator takes the traffic prefix before the current insertion position as input and outputs the corresponding burst length, with the LSTM hidden dimension set to 32, and the Softplus activation function [11] ensuring that the predicted length is non-negative. To maintain end-to-end trainability under discrete insertion operations, we adopt the Straight-Through Estimator (STE) [3] to round the predicted burst length.
The dynamic trigger generator simulates restricted-access conditions by randomly sampling insertion locations during training, with a loss function consisting of two components: a sequence-difference loss based on the fast Levenshtein distance ($D_{F}$) to enhance the distinguishability of post-defense traffic at the sequence-structure level, and a squared penalty term that constrains the total insertion length to not exceed the preset budget. These two terms are balanced by a trade-off coefficient ($\lambda=0.001$). On the Rimmer dataset, the dynamic trigger generator is optimized for approximately 1.7k training iterations using mini-batches of size 1024. During inference, the generator relies solely on the currently observable traffic prefixes for prediction without accessing the full traffic traces, thereby supporting defense deployment under online and restricted-access conditions. For the detailed defense algorithm and implementation details, please refer to Supplementary Materials E and F.
Deployment clarification. We distinguish between the training phase and the deployment phase of a defense mechanism. In the training phase, the defender selects a backdoor target label (e.g., a high-frequency website) and designs and verifies the defense offline in conjunction with the triggering perturbation. This phase is solely used for analyzing and constructing the defense mechanism itself and does not constitute an assumption about actual system deployment. Through this process, the association between the triggering perturbation and the target label in the attacker’s model is consolidated, thereby yielding a defense pattern that can be directly activated during subsequent deployment.
In the actual deployment phase, the defense neither needs to know nor infer the user’s access intent or differentiate among specific backdoor targets; instead, the defense uniformly injects the trigger perturbation into the protected web traffic. It is important to emphasize that TrapFlow does not rely on online learning after deployment; its defensive effect takes effect immediately upon deployment. Even if the attacker directly applies an existing attack model for inference without any retraining, the defended traffic still induces significant discriminative bias in the testing phase. If the attacker further retrains on newly collected data, the adaptation process instead reinforces the correlation between the trigger perturbations and the target labels, which in turn further degrades attack recovery.
During deployment, dynamic triggers are generated in an online manner, and the trigger generator relies solely on the currently observable traffic prefixes to predict the injection intensity, without accessing complete traffic traces or future information. For each protected site, all corresponding traffic traces are consistently injected with trigger perturbations, rather than sampling traffic, so as to prevent attackers from bypassing the defense through screening or statistical analysis. Trigger injection is implemented by inserting dummy incoming cells, which do not alter the original packet sending order; the timestamps of the inserted cells are generated via linear interpolation between adjacent original packets, thereby preserving the monotonicity of the time series and protocol consistency. This design is compatible with existing traffic stuffing and shaping mechanisms and is suitable for deployment in real systems such as Tor.
VI-B Closed-world Evaluation
@@
TrapFlow-s (heavy) 73.7 0.0 4.7 2.8 3.4 5.9 5.3 6.1
TrapFlow-d (heavy) 73.7 0.0 5.8 3.9 4.7 6.7 6.5 7.8
TrapFlow-d (light) 14.7 0.0 16.8 11.3 14.3 9.6 18.2 12.7
@@
In this subsection, we assess the effectiveness of six defense strategies across six WF attack methods. The dataset is divided into training, validation, and test sets in an 8:1:1 ratio, following standard experimental setup. To minimize potential bias from label selection, the accuracy of TrapFlow defense is calculated as the average over ten randomly selected target labels. We compare TrapFlow’s performance in both static (TrapFlow-s) and dynamic (TrapFlow-d) trigger patterns, with “light and “heavy” versions denoting the insertion of 4,000 and 20,000 incoming cells, respectively. Tab. I shows the average defense accuracy on the test set over ten random splits, with green arrows indicating the average reduction in attack accuracy achieved by TrapFlow compared to other defense methods for a certain attack.
@@
From Tab. I, we draw the following conclusions. ❶ Low overhead. TrapFlow maintains minimal time overhead (0.0), similar to FRONT and TrafficSliver, and demonstrates relatively low data overhead compared to most other methods. ❷ Superior defense performance. TrapFlow exhibits significant advantages in defense effectiveness. For instance, under the RF attack, the average accuracy reduction across TrapFlow’s variants exceeds 39.0, with TrapFlow-s (heavy) and TrapFlow-d (heavy) achieving accuracies of 4.7 and 5.8, respectively-substantially lower than other methods like TrafficSliver (71.0), WTF-PAD (98.3), and Palette (31.6). ❸ Broad applicability. TrapFlow’s “heavy” versions provide effective defense across all 6 WF attacks, with a maximum attack accuracy of 7.8. The “light” versions also achieve consistent performance, keeping the highest attack accuracy under 18.2. ❹ Light vs. heavy configurations. The “heavy” versions of TrapFlow-s and TrapFlow-d deliver stronger defense across all attacks but come with higher data overhead (73.7% for TrapFlow-d (heavy) vs. 14.7% for TrapFlow-d (light)). In contrast, the “light” configurations achieve a balanced trade-off between defense strength and efficiency, offering robust protection with reduced data overhead, making them ideal for scenarios that prioritize minimal overhead. ❺ Static vs. dynamic adaptability. While TrapFlow-s provide slightly stronger defense performance, they are less feasible for real-world deployment due to their reliance on fixed insertion locations, which reduces flexibility. Conversely, TrapFlow-d sacrifices a small amount of effectiveness but gains substantial adaptability by dynamically adjusting insertion locations in real time, making it more suitable for deployment in dynamic traffic environments. ❻ Defense advantages at comparable overhead. TrapFlow-s (heavy) demonstrates significant defensive superiority. For instance, in the DF attack, TrapFlow-s reduces accuracy to 2.8%, while FRONT’s accuracy remains at 86.2%, achieving an additional reduction of 83.4%. Compared to Tamaraw’s 11.1%, TrapFlow-s further lowers accuracy by 8.3%. In the RF attack, TrapFlow-s achieves an accuracy of 4.7%, while Palette reaches 31.6%, giving TrapFlow-s an extra reduction of 26.9%. These results indicate that TrapFlow-s offers stronger defense at similar data overheads. Compared with ALPaCA, TrapFlow-s (heavy) demonstrates significant defensive superiority across various attacks at similar overhead levels. TrapFlow-s (heavy) reduces the accuracy of DF attacks to 2.8%, whereas ALPaCA maintains an accuracy of 88.3%. In comparison, TrapFlow-s achieves an additional reduction of 85.5%. In RF attacks, TrapFlow-s (heavy) lowers the accuracy to 4.7%, while ALPaCA achieves 95.2% accuracy, resulting in an extra reduction of 90.5% by TrapFlow-s. As TrapFlow-d is practical for real-world deployment, our subsequent experiments mainly use the TrapFlow-d setup unless otherwise specified. We also show the comparison results of TrapFlow and backdoor-learning-based WF defenses in Supplementary Material G.
Robustness against classical feature-based WF attacks. We further evaluate the effectiveness of the proposed defense on classical feature-driven website fingerprinting attack methods, including CUMUL [51] and k-FP [21], which primarily rely on handcrafted statistical features rather than deep representation learning. Without defense, the attack accuracy of k-FP and CUMUL reaches 97.8% and 97.6%, respectively. Under the TrapFlow-s (heavy) defense, the attack accuracy of CUMUL and k-FP decreases to 5.1% and 5.4%, respectively, while the corresponding results are 6.0% and 6.5% for the TrapFlow-d (heavy) defense. Even with the lighter TrapFlow-d (light) defense, the attack accuracy is still effectively suppressed, reaching 14.3% for CUMUL and 17.2% for k-FP. Overall, the defense’s suppression effect across different feature-driven attack models is consistent with its performance on deep learning-based attack models, indicating that the method’s effectiveness does not depend on the specific model structure or feature extraction mechanism. A theoretical explanation of its model-independence is provided in Supplementary Material D.
VI-C Open-world Evaluation